# Generalizable Humanoid Manipulation with Improved 3D Diffusion Policies

### v1

@@ VI Capabilities
Data. We collect 10 demonstrations for each task. For Pick&Place task, each demonstration contains 10 trajectories of pick&place. In each demonstration, the object poses are randomized, limited in a region of $\times×10cm$. We do not collect data in a larger region, since we find that a larger task region simply requires more data [65]. Besides, collecting large-scale data is not feasible due to the usage of AVP.
Effectiveness. As shown in Table III, both iDP3 and DP achieve high success rates in the training environment with the training objects.
@@
Figure 8: View invariance of iDP3. We find that egocentric 3D representations are surprisingly view-invariant. Here DP corresponds to DP (✶R3M) in Table I, which is the strongest image-based baseline we have.
@@
VII Conclusions and Limitations
Conclusions. This work presents an imitation learning system that enables a full-sized humanoid robot to generalize practical manipulation skills to diverse real-world environments, trained with data collected solely in the lab. The key is the Improved 3D Diffusion Policy (iDP3), a new 3D visuomotor policy for general-purpose robots. Through extensive experiments, we demonstrate the impressive generalization capabilities of iDP3 in the real world.
Limitations. 1) Teleoperation with AVP is easy to set up but tiring for human teleoperators, making data scaling infeasible. 2) The depth sensor produces noisy point clouds, limiting the performance of iDP3. 3) Collecting fine-grained manipulation skills, such as turning a screw, is time-consuming due to teleoperation with AVP. 4) We avoided using the robot’s lower body, as maintaining balance is still challenging. In general, scaling up high-quality data is the main bottleneck. In the future, we hope to explore how to scale up the training of 3D visuomotor policies with more high-quality data.
Acknowledgments
We would like to thank Jie Gu, Bin Zhou, and Yusheng Cai from Fourier Intelligence for hardware support, Yuxiang Gao for help in teleoperation, Shuo Hu for help in the 3D print of the camera mount, Zeqian Bao, Renzhi Tao, and Jiayan Gu for helpful discussions. Besides, we would like to thank Chen Wang, Yunzhi Zhang, Zizhang Li, and Haoyu Xiong from Stanford University for their insightful discussions.

### v2

@@ IV-D Capabilities
[…]
Effectiveness. As shown in Table IV, both iDP3 and DP achieve high success rates in the training environment with the training objects.
@@
Figure 8: View invariance of iDP3. We find that egocentric 3D representations are surprisingly view-invariant. Here DP corresponds to DP (✶R3M) in Table II, which is the strongest image-based baseline we have.
@@
V Conclusions and Limitations
Conclusions. This work presents a real-world imitation learning system that enables a full-sized humanoid robot to generalize practical manipulation skills to diverse real-world environments, trained with data collected solely in one single scene. With more than 2000 rigorous evaluation trials, we present an improved 3D Diffusion Policy, that can learn robustly from human data and perform effectively on our humanoid robot. The results that our humanoid robot can perform autonomous manipulation skills in diverse real-world scenes show the potential of using 3D visuomotor policies in real-world manipulation tasks with data efficiency.
Limitations. 1) Teleoperation with Apple Vision Pro is easy to set up, but it is tiring for human teleoperators, making imitation data hard to scale up within the research lab. 2) The depth sensor still produces noisy and inaccurate point clouds, limiting the performance of iDP3. 3) Collecting fine-grained manipulation skills, such as turning a screw, is time-consuming due to teleoperation with AVP; systems like Aloha [18] are easier to collect dexterous manipulation tasks at this stage. 4) We avoided using the robot’s lower body, as maintaining balance is still challenging due to the hardware constraints brought by current humanoid robots. In general, scaling up high-quality manipulation data is the main bottleneck. In the future, we hope to explore how to scale up the training of 3D visuomotor policies with more high-quality data and how to employ our 3D visuomotor policy learning pipeline to humanoid robots with whole-body control.